The measures of two angles are in the ratio 4:5. The larger of the two angles is 60 less than twice the smaller angle. Find the measure of each angle. Are the angles complementary? Supplementary?
Answer by lwsshak3(11628)

The measures of two angles are in the ratio 4:5. The larger of the two angles is 60 less than twice the smaller angle. Find the measure of each angle. Are the angles complementary? Supplementary?
***
let x=smaller angle
y=larger angle
x/y=4/5
x=(4/5)y
..
y=2x-60
y=(8/5)y-60
5y=8y-300
3y=300
y=100
x=(4/5)y=80
smaller angle=80
larger angle=100
angles are supplementary